Ms. Riley deposited some money in an account paying 5% simple interest and twice that amount in an account paying 6% simple interest. If the total interest from the two accounts is 765$ for 1 year, how much was deposited into each account?
Answer by mananth(16949)

Part I 5.00% per annum ------------- x
Part II 6.00% per annum ------------ y=2x
total investment 30000
Interest----- $765.00

5.00% x + 6.00% * 2 x = $765.00
0.05 x + 0.06 * 2 x = $765.00
0.05 x + 0.12 x = $765.00
0.17 x = $765.00
/0.17
x=$4,500.00--- Part I
2x=$9,000.00 Part II